OpenStudy (anonymous):

solve the right triangles with the given parts. round results. a=932,c=1240

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

assuming this is a right triangle, you use the Pythagorean theorem a^2 + b^2 = c^2 932^2 + b^2 = 1240^2 868624 + b^2 = 1537600 now solve for b

jimthompson5910 (jim_thompson5910):

not sure how you're getting that

OpenStudy (anonymous):

i mean b^2=668976

OpenStudy (anonymous):

is ubtracted 868624 fromb oth sides and that leaves me with b^2=668976 is this correct sof ar?

jimthompson5910 (jim_thompson5910):

yes it is, what's next?

OpenStudy (anonymous):

divide

jimthompson5910 (jim_thompson5910):

no, what undoes squaring?

OpenStudy (anonymous):

square root?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

so you have to take the square root of both sides

OpenStudy (anonymous):

i get 817.9 rounds to 818

OpenStudy (anonymous):

b=818

jimthompson5910 (jim_thompson5910):

if you're rounding to the nearest whole number, then b = 818 is correct

OpenStudy (anonymous):

i mean tan^-1 818/932=41.3

OpenStudy (anonymous):

then take 90-41.3=48.7 if i rounded correctly

OpenStudy (anonymous):

to the correct significant digit

jimthompson5910 (jim_thompson5910):

not sure what you're doing now

jimthompson5910 (jim_thompson5910):

you're looking for the angles right?

OpenStudy (anonymous):

right triangles with given parts but i think i'm su pposed to find the other angles a s well

OpenStudy (anonymous):

i get A= 48.73 and B=41.27 correct?

jimthompson5910 (jim_thompson5910):

that is correct, I'm getting the same thing as well pretty much